solve y(1+2xycosx^2-2xy)dx+(sinx^2-x^2)dy=0


1
Expert's answer
2021-02-24T06:42:24-0500

NOTE: I asked for cosx^2 is cos(x2) or cos2x at the work start but no answer.

Solution

y(1+2xycos2x-2xy)dx + (sin2x-x2)dy = 0

Dividing by (sin2x-x2)dx we will get

dy/dx = -y(1+2xycos2x-2xy)/(sin2x-x2) =

           = y(2xysin2x - 1)/(sin2x-x2) = -y/(sin2x-x2)+ 2x y2 sin2x /(sin2x-x2)

So y’+ y/(sin2x-x2) = [2x sin2x /(sin2x-x2)] y2

It is Bernoulli differential equation

y’ + p(x)y = q(x)y2  

with p(x) = (sin2x-x2)-1 , q(x) = 2x sin2x /(sin2x-x2

Divide this by y2 .

y-2y’ + p(x)y-1 = q(x)   

The substitution and derivative that we’ll need here is

v=y−1, v′=−y−2y′

With this substitution the differential equation becomes

v’ – p(x)v = -q(x)

Let m(x) = exp(-∫p(x) dx) = exp(-∫(sin2x-x2)-1  dx)

Multiplying last equation m(x) we’ll get

(m(x)v)’ = -m(x) q(x)

Integrating this ∫ (m(x)v)’dx = -∫m(x) q(x)dx and so

v = 1/y = -[m(x)]-1∫ m(x) q(x)dx - C

And y = -{[m(x)]-1∫ m(x) q(x)dx + C}-1  

Finally

y = -{[ exp(-∫(sin2x-x2)-1  dx)]-1∫ exp(-∫(sin2x-x2)-1  dx)*2x sin2x /(sin2x-x2) dx + C}-1

or

"y(x)=\ -\{ \ [ exp(-\int \frac{dx}{sin^2x-x^2} ) \ ]^{-1} 
\int exp(-\int \frac{dx}{sin^2x-x^2} ) \frac{2x sin^2x}{sin^2x-x^2} dx +C \}^{-1}"